• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 16:00:24 on December11,2008

    Entry number 252998

    keyword=Day Shift Summary

    Summary Day Shift 11-Dec-08

    Ole Hansen (SL)
    Wolfgang Korsch (TO)
    Elena Long (3rd)


    Summary:

    Smooth data taking with good beam. Beam was down for about 45 minutes
    in the morning due to a magnet trip. Compton polarimeter continues to
    have high backgrounds, but expert managed to bring it back to a state
    where the data "are useful".

    Run summary: (all runs on 3He in "pi+ transverse" kinematics)


    run# #events HAPPEX run# Current Charge Comment
    (uA) (mC)
    4394 1M 30551 12 ? (incomplete end-of-
    run record)
    4396 1M 30551 12 38.22
    4397 1M 30552 12 39.40
    4398 294k 30552 12 10.13
    4399 1M 30552 12 39.46
    4400 1M 30552 12 39.66
    4401 1M 30553 12 39.34
    4402 1M 30553 12 (in progress at
    end of shift)